OpenStudy (anonymous):

There are 10 vehicles in a parking lot: 3 SUVs and 7 trucks. What is the probability that of any 7 randomly chosen vehicles, exactly 1 is an SUV?

OpenStudy (mathmate):

S=SUV T=truck 7 possible outcomes for exactly 1 SUV STTTTTT TSTTTTT TTSTTTT TTTSTTT TTTTSTT TTTTTST TTTTTTS Each of these outcomes has a probability of \[\frac{3}{10}\frac{7}{9}\frac{6}{8}\frac{5}{7}\frac{4}{6}\frac{3}{5}\frac{2}{4}=\frac{1}{10}\] So with all 7 possible outcomes, probability = 7/10

OpenStudy (anonymous):

Can someone please explain how to find the possible outcome f having 7 vehicles in a parking lot.

OpenStudy (mathmate):

@melisamangeee @anjelina24dtl On second thought, I would do it differently using multinomial arrangements. X=Number of ways of arranging 7 vehicles, out of which 1 is an SUV = 7!/(6!1!) =7 N=Number of ways of arranging 10 vehicles, 3 of which are SUV's =10!/(7!3!) =120 Probability of choosing 1 SUV among 7 total vehicles out of 7 trucks and 3 SUV's = 7/120